OpenStudy (anonymous):

The average annual rainfall for a town is 43.2 inches. What is the average monthly rainfall? The average monthly rainfall for the previous 9 months was 4 inches. Did the town exceed its average annual rainfall? If so, by how much?

OpenStudy (anonymous):

The average is 3.6

OpenStudy (anonymous):

thank you!! how did you get the answer?

OpenStudy (anonymous):

yes it did exceed the average rain fall

OpenStudy (anonymous):

you divide 43.2 divided by 12

OpenStudy (texaschic101):

average annual rainfall is 43.2 inches.....there are 12 months in a year, so divide by 12 to get the average monthly rainfall. 43.2/12 = 3.6 per month

OpenStudy (anonymous):

it exceeded by 10.8 inches

OpenStudy (texaschic101):

9 months at 4 inches per month = (9 * 4) = 36 inches I am a little confused here....are we supposed to compare what was made for the 9 months to the average yearly rainfall ?

OpenStudy (anonymous):

what i did is 3.6 x 3 because of the 3 months missing, then i got 10.8 and that is it

OpenStudy (texaschic101):

my question is : are we supposed to account for the 3 missing months

OpenStudy (anonymous):

yes because it says did it exceed and by how much so you count the whole year